# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

### v1

@@ 1.3. Organization of the rest of the paper
2. Notation
Let ${\mathbb{N}}$ be the set of all strictly positive integers and ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$. Fix $d\in{\mathbb{N}}$ and $T\in(0,\infty)$. Let ${\mathbb{R}}_{+}=[0,\infty)$. Let $E={\mathbb{R}}^{d}$ with cemetery state $\Delta\not\in E$ Given $(s,x)\in{\mathbb{R}}_{+}\times D({\mathbb{R}}_{+},E)$ and $e\in E$, define
@@ 2. Notation
Given a topological space $S$, we denote by $\mathcal{B}(S)$ its Borel $\sigma$-field, by $\mathscr{P}(S)$ the set of all probability measures on $\mathcal{B}(S)$, and by $B(S)$ the set of all bounded Borel-measurable functions from $S$ to ${\mathbb{R}}$. Moreover, we write $C(S)$, $\mathrm{USC}(S)$, and $\mathrm{LSC}(S)$ for the sets of all continuous, upper semicontinuous, and lower semicontinuous functions from $S$ to ${\mathbb{R}}$, resp. Given another topological space $\tilde{S}$, we write $C(S,\tilde{S})$ for the set of all continuous functions from $S$ to $\tilde{S}$.
@@ 3.1. The canonical path space
Let $\Omega=D({\mathbb{R}}_{+},E)$. The canonical process $X=(X_{t})_{t\geq 0}$ on $\Omega$ is defined by $X_{t}(\omega)=\omega(t)$ for each $(t,{\omega})\in{\mathbb{R}}_{+}\times\Omega$. The random measure associated to the jumps of $X$ is denoted by $\mu^{X}$, i.e.,
$\mu^{X}({\mathrm{d}}t\,{\mathrm{d}}e)=\sum_{r}{\mathbf{1}}_{\{\Delta X_{r}\neq 0\}}\,\delta_{(r,X_{r})}({\mathrm{d}}t\,{\mathrm{d}}e).$
Here, $\Delta X_{r}=X_{r}-X_{r-}$. Let ${\mathbb{F}}^{0}=\{{\mathcal{F}}^{0}_{t}\}_{t\geq 0}$ be the filtration generated by $X$ and equip $\Omega$ with the $\sigma$-field ${\mathcal{F}}^{0}:=\vee_{t\geq 0}{\mathcal{F}}^{0}_{t}$. We also use the right-continuous filtration ${\mathbb{F}}=\{{\mathcal{F}}_{t}\}_{t\geq 0}$ defined by ${\mathcal{F}}_{t}:=\cap_{{\varepsilon}>0}{\mathcal{F}}^{0}_{t+{\varepsilon}}$ and the $\sigma$-field ${\mathcal{F}}:=\vee_{t\geq 0}{\mathcal{F}}_{t}$.
We consider non-empty subsets of +×${\mathbb{R}}_{+}\times\Omega$ to be equipped with the pseudo-metric
@@ 3.2. The canonical space of controlled marked point processes
[…]
@@
(i) $(t_{n},e_{n},\alpha_{n})\in\left((0,\infty)\times E\times\mathcal{A}\right)\cup\{(\infty,\Delta,\Delta^{\prime})\}$.
@@
For each $n\in{\mathbb{N}}_{0}$, define canonical mappings ${T}_{n}:\check{\Omega}\to(0,\infty]$, ${E}_{n}:\check{\Omega}\to E\cup\{\Delta\}$, and $\mathcal{A}_{n}:\check{\Omega}\to\mathcal{A}\cup\{\Delta^{\prime}\}$ by
${T}_{n}(\check{\omega})=t_{n},\quad{E}_{n}(\check{\omega})=e_{n},\quad\mathcal{A}_{n}(\check{\omega})=\alpha_{n},\qquad\check{\omega}=(t_{n},e_{n},\alpha_{n})_{n\in{\mathbb{N}}_{0}}\in\check{\Omega}.$
@@
let $\check{{\mathbb{F}}}=\{\check{{\mathcal{F}}}_{t}\}_{t\geq 0}$ be the filtration [^8] This is a raw filtration. It is emphasized in [Jacobsen] that completed filtrations should not be used. generated by $\check{\mu}$, and equip $\check{\Omega}$ with the $\sigma$-field ˇ$\check{{\mathcal{F}}}:=\vee_{t\geq 0}\check{{\mathcal{F}}}_{t}$.
@@
The filtration $\check{{\mathbb{F}}}$ is right-continuous (Theorem T25, p. 304, in [BremaudBook]).
@@
$\check{\Omega}_{n}$ :=$:=\{\check{\omega}_{n}\in(({\mathbb{R}}_{+}\times E\times\mathcal{A})\cup\{(\infty,\Delta,\Delta^{\prime})\})^{n}:$
@@
with corresponding filtration $\check{{\mathbb{F}}}_{n}=\{\check{{\mathcal{F}}}_{n,t}\}_{t\geq 0}\}$ and $\sigma$-field ˇ$\check{{\mathcal{F}}}_{n,\infty}:=\vee_{t\geq 0}\check{{\mathcal{F}}}_{n,t}$. With slight abuse of notation, we write $T_{j}$, $E_{j}$, and $\mathcal{A}_{j}$ for the canonical mappings on $\check{\Omega}_{n}$ and do not indicate their domains (i.e., $\check{\Omega}_{n}$ vs. $\check{\Omega}$) if there is no danger of confusion.
@@ 4.1. The data
$f:{\mathbb{R}}_{+}\times\Omega\times A\to E,\quad\lambda:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}_{+},\quad Q:{\mathbb{R}}_{+}\times\Omega\times A\times\mathcal{B}(E)\to{\mathbb{R}}_{+},$
which are the characteristics for our PDP. Next, consider mappings
:+×$\ell:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}_{+}\qquad\text{and}\qquad h:\Omega\to{\mathbb{R}}_{+},$
@@
(i) For a.e. $t\in{\mathbb{R}}_{+}$, the map $(x,a)\mapsto(f,\lambda,\ell)(t,x,a)$, $\Omega\times A\to E\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}$, is continuous, where $\Omega$ is considered here to be equipped with the seminorm $x\mapsto\sup_{s\leq t}\left|{x(s\wedge t)}\right|$.
(ii) For every $(x,a)\in\Omega\times A$, the map $t\mapsto(f,\lambda,\ell)(t,x,a)$, ${\mathbb{R}}_{+}\to E\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}$, is Borel-measurable.
(iii) There are constants $C_{f}$, $C_{\lambda}\geq 0$, $L_{f}\geq 0$, such that, for every $(s,x)\in{\mathbb{R}}_{+}\times\Omega$,
@@
(i) For every $B\in\mathcal{B}(E)$, $(s,x,a)\mapsto Q(s,x,a,B)$ is Borel-measurable and non-anticipating, i.e., $Q(s,x,a,B)=Q(s,x(\cdot\wedge s),a,B)$.101010 By Theorem 97, 147-IV, in [DMprobPot], the maps $(s,x)\mapsto Q(s,x,a,B)$, $(a,B)\in A\times\mathcal{B}(E)$, are ${\mathbb{F}}^{0}$-optional.
(ii) For every $s\in{\mathbb{R}}_{+}$, $x\in\Omega$, and $a\in A$, $Q(s,x,a,\cdot)\in\mathscr{P}(E)$.
@@
$\left|{\psi(s,x)-\psi(s,\tilde{x})}\right|\leq L_{\psi}{\color[rgb]{0,0,0}\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|}\text{ for all $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$}$
@@
$\qquad\leq L_{\psi}\,L_{Q}\,\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|\text{ for all $(s,x,\tilde{x},e)\in[0,T]\times\Omega\times\Omega\times E$.}$
@@
By Proposition 7.26 in [BS], $Q$ is a Borel-measurable stochastic kernel on $E$ given +×${\mathbb{R}}_{+}\times\Omega\times A$.
@@ 4.2. The flow and related notation
Establishing existence and uniqueness for the initial value problem (4.1) is standard given Assumption 4.1. Note that $\phi^{s,x,\alpha}\in\Omega$ and $\phi^{s,x,\alpha}|_{[s,\infty)}\in C([s,\infty),E)$.
@@
The map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$.
@@
$Q(t,{\mathrm{d}}e|s,x,\alpha):=Q^{s,x,\alpha}(t,{\mathrm{d}}e),\text{ etc.}$.
@@
Following Lemma 4.4 and the arguments in the proofs in section B.1,111111 E.g., assume first that $\lambda=\lambda(t,x,a)$ is of the form $\lambda_{1}(t)\,\lambda_{2}(x(\cdot\wedge t))\,\lambda_{3}(a)$, where $\lambda_{1}$ is (Borel) measurable and $\lambda_{2}$ as well as $\lambda_{3}$ are continuous. Then, by Lemma 4.4, $(s,x,\alpha,t)\mapsto\lambda(t,\phi^{s,x,\alpha},\alpha(t))$ is clearly $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})\otimes\mathcal{B}({\mathbb{R}}_{+})$-measurable. It remains to apply a monotone-class argument. one can deduce that the mappings
@@
$(s,x,\alpha,t)\mapsto Q^{s,x,\alpha}(t,B),{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},B\in\mathcal{B}(E),$
@@
4.3. Path-dependent controlled PDPs on $\check{\Omega}$
Let $(s,x)\in(0,\infty)\times\Omega$. Define the process $X^{s,x}:{\mathbb{R}}_{+}\times\check{\Omega}\to E\cup\{\Delta\}$ recursively by

### v2

@@ 1.5. Organization of the rest of the paper
2. Notation and preliminaries
Let $\mathbb{N}$} be the set of all strictly positive integers and ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$. Fix $d\in{\mathbb{N}}$ and $T\in(0,\infty)$. Let $\Delta\not\in{\mathbb{R}}^{d}$ be a cemetery state. Given $(s,x)\in{\mathbb{R}}_{+}\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$ and $e\in{\mathbb{R}}^{d}$, define
@@ 2. Notation and preliminaries
Given sets $S$ and $\tilde{S}$, we denote by $\tilde{S}^{S}$ the set of all functions from $S$ to $\tilde{S}$.
Given a topological space $S$, we denote by $\mathcal{B}(S)$ its Borel $\sigma$-field, by $\mathscr{P}(S)$ the set of all probability measures on $\mathcal{B}(S)$, and by $B(S)$ the set of all bounded Borel-measurable functions from $S$ to $\mathbb{R}$}. Moreover, we write $C(S)$, $C_{b}(S)$, $\mathrm{USC}(S)$, $\mathrm{BUSC}(S)$, $\mathrm{LSC}(S)$, and $\mathrm{BLSC}(S)$, resp., for the sets of all continuous, bounded continuous, upper semicontinuous, bounded upper semicontinuous, lower semicontinuous, and bounded lower semicontinuous functions from $S$ to $\mathbb{R}$}, resp. Given another topological space $\tilde{S}$, we write C(S,S~)C($\tilde{S}$) for the set of all continuous functions from $S$ to $\tilde{S}$. If $S$ is a Borel space, then we write $\mathrm{BLSA}(S)$ for the set of all bounded lower semi-analytic functions from $S$ to $\mathbb{R}$} (for a detailed treatment of analytic sets and lower semi-analytic functions, we refer the reader to sections 7.1 and 7.2 in [BS]).
Let us also recall the notion of universal measurability: A subset $\underline{S}$ of a Borel space $S$ is universally measurable if, for every ${{\mathbb{P}}}\in\mathscr{P}(S)$, the set $\underline{S}$ belongs to the completion of $\mathcal{B}(S)$ with respect to $\mathbb{P}$}} (see Definition 7.18 of [BS]).
@@
Occasionally, we use $\lVert{x}\rVert_{s}:=\sup_{0\leq t\leq s}\left|{x(t)}\right|$ given $s\in[0,T]$ and $x\in D({\mathbb{R}}_{+},E)$. Moreover, given a bounded real-valued function $\varphi$, we denote its sup norm by $\lVert{\varphi}\rVert_{\infty}$.
@@ 3.1. The canonical path space
Let $\Omega=D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$. The canonical process $X=(X_{t})_{t\geq 0}$ on $\Omega$ is defined by $X_{t}(\omega)=\omega(t)$ for each $(t,{\omega})\in{\mathbb{R}}_{+}\times\Omega$. Let ${\mathbb{F}}^{0}=\{{\mathcal{F}}^{0}_{t}\}_{t\geq 0}$ be the filtration generated by $X$ and let $\Omega$ be equipped with the $\sigma$-field ${\mathcal{F}}^{0}:=\vee_{t\geq 0}{\mathcal{F}}^{0}_{t}$.
We consider non-empty subsets of ${\mathbb{R}}_{+}\times\Omega$ to be equipped with the pseudo-metric
@@ 3.2. The canonical space of controlled marked point processes
$\mathcal{A}:=\{\alpha:{\mathbb{R}}_{+}\to A\text{ Borel-measurable}\}$
@@
(i) $(t_{n},e_{n},\alpha_{n})\in\left((0,\infty)\times{\mathbb{R}}^{d}\times\mathcal{A}\right)\cup\{(\infty,\Delta,\Delta^{\prime})\}$.
@@
For each $n\in{\mathbb{N}}_{0}$, we define canonical mappings ${T}_{n}:\check{\Omega}\to(0,\infty]$, ${E}_{n}:\check{\Omega}\to{\mathbb{R}}^{d}\cup\{\Delta\}$, and $\mathcal{A}_{n}:\check{\Omega}\to\mathcal{A}\cup\{\Delta^{\prime}\}$ by
[…]
and we also write $\mathcal{A}_{n}(t)$ instead of $\check{\omega}\mapsto[\mathcal{A}_{n}(\check{\omega})](t)$, $\check{\Omega}\to A$, $t\in{\mathbb{R}}_{+}$, i.e.,
$[\mathcal{A}_{n}(t)](\check{\omega})=\alpha_{n}(t),\qquad\check{\omega}=(t_{j},e_{j},\alpha_{j})_{j\in{\mathbb{N}}_{0}}\in\check{\Omega}.$
@@
let $\check{{\mathbb{F}}}=\{\check{{\mathcal{F}}}_{t}\}_{t\geq 0}$ be the filtration generated by $\check{\mu}$, and equip $\check{\Omega}$ with the $\sigma$-field $\check{{\mathcal{F}}}:=\vee_{t\geq 0}\check{{\mathcal{F}}}_{t}$.
@@
The filtration $\check{{\mathbb{F}}}$ is right-continuous (Theorem T25, p. 304, in [BremaudBook]) and, by definition, is a raw filtration. It is emphasized in Remark 4.2.1 of [Jacobsen] that it is “essential” not to use completions of $\check{{\mathbb{F}}}$ in the context of the construction of marked point processes.
@@
$\check{\Omega}_{n}$ $:=\{\check{\omega}_{n}\in(({\mathbb{R}}_{+}\times E\times\mathcal{A})\cup\{(\infty,\Delta,\Delta^{\prime})\})^{n}:$
@@
with corresponding filtration $\check{{\mathbb{F}}}_{n}=\{\check{{\mathcal{F}}}_{n,t}\}_{t\geq 0}\}$ and $\sigma$-field $\check{{\mathcal{F}}}_{n,\infty}:=\vee_{t\geq 0}\check{{\mathcal{F}}}_{n,t}$. With slight abuse of notation, we write $T_{j}$, $E_{j}$, and $\mathcal{A}_{j}$ for the canonical mappings on $\check{\Omega}_{n}$ and do not indicate their domains (i.e., $\check{\Omega}_{n}$ vs. $\check{\Omega}$) if there is no danger of confusion.
@@ 4.1. The data
$f:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}^{d},\quad\lambda:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}_{+},\quad Q:{\mathbb{R}}_{+}\times\Omega\times A\times\mathcal{B}({\mathbb{R}}^{d})\to{\mathbb{R}}_{+},$
which are the characteristics for our PDP, i.e., $f$ specifes its deterministic flow, $\lambda$ the distribution of its jump times, and $Q$ the distribution of its post-jump locations. Next, consider mappings
$\ell:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}_{+}\qquad\text{and}\qquad h:\Omega\to{\mathbb{R}}_{+},$
@@
(i) For a.e. $t\in{\mathbb{R}}_{+}$, the map $(x,a)\mapsto(f,\lambda,\ell)(t,x,a)$, $\Omega\times A\to{\mathbb{R}}^{d}\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}$, is continuous, where $\Omega$ is considered here to be equipped with the seminorm $x\mapsto\sup_{s\leq t}\left|{x(s\wedge t)}\right|$.
(ii) For every $(x,a)\in\Omega\times A$, the map $t\mapsto(f,\lambda,\ell)(t,x,a)$, ${\mathbb{R}}_{+}\to{\mathbb{R}}^{d}\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}$, is Borel-measurable.
(iii) There are constants $C_{f}$, $C_{\lambda}\geq 0$, such that, for every $(s,x)\in{\mathbb{R}}_{+}\times\Omega$,
@@
(i) For every $B\in\mathcal{B}({\mathbb{R}}^{d})$, $(s,x,a)\mapsto Q(s,x,a,B)$ is Borel-measurable and non-anticipating, i.e., $Q(s,x,a,B)=Q(s,x(\cdot\wedge s),a,B)$.
(ii) For every $s\in{\mathbb{R}}_{+}$, $x\in\Omega$, and $a\in A$, $Q(s,x,a,\cdot)\in\mathscr{P}({\mathbb{R}}^{d})$.
@@
(4.1) $\left|{\psi(s,x)-\psi(s,\tilde{x})}\right|\leq L_{\psi}{\color[rgb]{0,0,0}\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|}\text{ for all $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$}$
@@
$\qquad\leq L_{\psi}\,L_{Q}\,\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|\text{ for all $(s,x,\tilde{x},e)\in[0,T]\times\Omega\times\Omega\times{\mathbb{R}}^{d}$.}$
(v) For each $s\in{\mathbb{R}}_{+}$ and $\eta\in C_{b}({\mathbb{R}}^{d})$, the map $(x,a)\mapsto\int_{{\mathbb{R}}^{d}}\eta(e)\,Q(s,x,a,{\mathrm{d}}e)$, $\Omega\times A\to{\mathbb{R}}$, is continuous.
@@
By Proposition 7.26 in [BS], $Q$ is a Borel-measurable stochastic kernel on ${\mathbb{R}}^{d}$ given ${\mathbb{R}}_{+}\times\Omega\times A$. By Theorem 97 (b), 147-IV, in [DMprobPot], the maps $(s,x)\mapsto Q(s,x,a,B)$, $(a,B)\in A\times\mathcal{B}({\mathbb{R}}^{d})$, are ${\mathbb{F}}^{0}$-optional.
We provide now two elementary examples concerning Assumption 4.2 (iv).
Example $4.$
Let $d=1$.
(i) Assume that
$Q(s,x,a,{\mathrm{d}}e)=\frac{1}{2}[\delta_{x(s)-1}({\mathrm{d}}e)+\delta_{x(s)+1}({\mathrm{d}}e)].$
Consider a function $\psi\in C_{b}([0,T]\times\Omega{\color[rgb]{0,0,0})}$ and a constant $L_{\psi}\geq 0$ such that (4.1) holds. Then, for every $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$, we have
$\Big|\int_{{\mathbb{R}}^{d}}\psi(s,x\otimes_{s}e)\,Q(s,x,a,{\mathrm{d}}e)-\int_{{\mathbb{R}}^{d}}\psi(s,x\otimes_{s}e)\,Q(s,\tilde{x},a,{\mathrm{d}}e)\Big|$
$=\frac{1}{2}\Big|\psi(s,x\otimes_{s}(x(s)-1))+\psi(s,x\otimes_{s}(x(s)+1))$
$\qquad\qquad-\psi(s,x\otimes_{s}(\tilde{x}(s)-1))-\psi(s,x\otimes_{s}(\tilde{x}(s)+1))\Big|$
$\color[rgb]{0,0,0}\leq L_{\psi}\left|{x(s)-\tilde{x}(s)}\right|$
$\leq L_{\psi}\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|,$
i.e., Assumption 4.2 (iv) holds.
(ii) Assume that
$Q(s,x,a,{\mathrm{d}}e)$ $=\begin{cases}\frac{1}{2}[\delta_{x(0)-1}({\mathrm{d}}e)+\delta_{x(0)+1}({\mathrm{d}}e)]&\text{ if $x(s)\not\in\{x(0)-1,x(0)+1\}$,}\\ \delta_{x(0)}({\mathrm{d}}e)&\text{ if $x(s)\in\{x(0)-1,x(0)+1\}$.}\end{cases}$
Similarly as in part (i), one can show that Assumption 4.2 (iv) holds. Also note that Assumption 4.2 (iii) holds as well.
@@ 4.2. The flow and related notation
Establishing existence and uniqueness for the initial value problem (4.2) is standard given Assumption 4.1 (see, e.g., Theorem 16.3.11 in [CohenElliott] for a fairly general result, which covers our case). Note that $\phi^{s,x,\alpha}\in\Omega$ and $\phi^{s,x,\alpha}|_{[s,\infty)}\in C([s,\infty),{\mathbb{R}}^{d})$.
@@
The map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}$, ${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$.
@@
$Q(t,{\mathrm{d}}e|s,x,\alpha):=Q^{s,x,\alpha}(t,{\mathrm{d}}e),\text{ etc.}$
@@
Following Lemma 4.5 and the arguments in the proofs in section B.1, one can deduce that the mappings
@@
$(s,x,\alpha,t)\mapsto Q^{s,x,\alpha}(t,B),{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},B\in\mathcal{B}({\mathbb{R}}^{d}),$
@@
To this end, one can start by assuming first that $\lambda=\lambda(t,x,a)$ is of the form $\lambda_{1}(t)\,\lambda_{2}(x(\cdot\wedge t))\,\lambda_{3}(a)$, where $\lambda_{1}$ is (Borel) measurable and $\lambda_{2}$ as well as $\lambda_{3}$ are continuous. Then, by Lemma 4.5, $(s,x,\alpha,t)\mapsto\lambda(t,\phi^{s,x,\alpha},\alpha(t))$ is clearly $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})\otimes\mathcal{B}({\mathbb{R}}_{+})$-measurable. It remains to apply a monotone-class argument.
4.3. The continuous-time optimal control problem
First, let us define recursively the “canonical” piecewise deterministic process $X^{s,x}:{\mathbb{R}}_{+}\times\check{\Omega}\to{\mathbb{R}}^{d}\cup\{\Delta\}$ starting at $(s,x)\in(0,\infty)\times\Omega$ by